# Is the $3S$-$2D$ mixing strong for the charmonia $\psi(4040)$ and $\psi(4160)$?

### v1

@@ Abstract
In this work, we revisit the $3S$-$2D$ mixing scheme for the charmonia $\psi(4040)$ and $\psi(4160)$. We introduce a coupled-channel mechanism—distinct from the tensor-force contribution in potential models, which alone is insufficient to induce significant mixing—to describe the mixing between these states. Our analysis yields mixing angles of $\theta_{1}=7^{\circ}$ and $\theta_{2}=10^{\circ}$, inconsistent with the larger angle inferred from experimental data, such as the di-lectronic widths of the $\psi(4040)$ and $\psi(4160)$. We discuss possible origins of this discrepancy and emphasize the need for future experiments to resolve it. Precise measurements of the resonance parameters and di-lectronic decay widths, via both inclusive and exclusive processes, will be crucial in clarifying this issue.
@@ I Introduction
Although fifty years have passed since the discovery of the $J/\psi$, the properties of charmonium states above the $D\bar{D}$ threshold remain poorly understood, presenting a more complex picture due to insufficient experimental data and a lack of comprehensive theoretical understanding. Among these higher-mass states, the $\psi(4040)$ and $\psi(4160)$ were first observed in $e^{+}e^{-}$ annihilation by the DASP Collaboration in 1978 [1]. Both states are now well-established charmonium resonances and are listed in the Review of Particle Physics by the Particle Data Group (PDG). This inclusion might suggest that their properties are sufficiently understood—yet in reality, the situation remains unclear, as we will discuss below.
$S$-$D$ mixing is a universal phenomenon in charmonium physics. A classic example is the $2S$-$1D$ mixing scheme proposed by Rosner for $\psi(3770)$ and $\psi(3686)$ [2], which successfully explains the relevant experimental data. Recently, the Lanzhou group suggested $4S$-$3D$ mixing as a mechanism to interpret the charmonium-like states $Y(4220)$ and $Y(4380)$ [3]. Similarly, for the well-established charmonia $\psi(4040)$ and $\psi(4160)$, a $3S$-$2D$ mixing scheme has also been proposed, as discussed in Refs. [4, 5, 6, 7, 8, 9, 10, 11]. The mixing angles can be roughly obtained by fitting either the measured decay widths of the relevant states or the ratio between them. In Table 1, we summarize the current understanding of $3S$-$2D$ mixing, including the inferred mixing angles and the corresponding approaches employed to extract them.
Table 1: The obtained $3S$-$2D$ mixing angle for the $\psi(4040)$ and $\psi(4160)$.
@@
Although various methods are employed to determine the mixing angles, their results indicate that a significant $3S$-$2D$ mixing effect exist. Understanding the nature of the mixing angle is essential, as it has an influence on the assignment of higher vector charmonium states, their properties of mass, production, and decay. Therefore, given the significant $3S$-$2D$ mixing effect observed in the charmonium states $\psi(4040)$ and $\psi(4160)$, it becomes crucial to identify the underlying dynamical mechanisms responsible for generating such strong mixing.
In this work, we focus on this issue. In the traditional potential model, the tensor term in the effective potential contributes to $S$-$D$ mixing in charmonium, but its effect is too weak to produce a large mixing angle. This necessitates exploring alternative mechanisms. With advances in understanding unquenched effects in hadron spectroscopy, higher-mass states can no longer be accurately described using the quenched approximation. Among these effects, the coupled-channel mechanism has emerged as particularly significant, signaling that hadron spectroscopy is entering a high-precision era [15, 16, 17, 18, 19]. Leveraging these developments, we investigate whether coupled-channel effects could explain the large $3S$-$2D$ mixing angle reported in earlier studies. However, we find a notable discrepancy between our calculations and prior results: our analysis yields mixing angles of $\theta_{1}=7^{\circ}$ and $\theta_{2}=10^{\circ}$, differing substantially from previous values.
This discrepancy cannot be overlooked. To resolve it, we scrutinize earlier methodologies and identify a key sensitivity: the mixing angle depends critically on the measured di-electronic widths of the $\psi(4040)$ and $\psi(4160)$. Unfortunately, existing experimental data are insufficient for a precise extraction of the mixing angle. Our findings underscore the urgent need for higher-precision measurements to clarify the $S$-$D$ mixing problem in the $\psi(4040)$ and $\psi(4160)$.
@@
II $3S$-$2D$ Mixing scheme for $\psi(3^{3}S_{1})$ and $\psi(2^{3}D_{1})$
As mentioned above, charmonia $\psi(3770)$ and $\psi(3686)$ can be well understood by incorporating $2S$-$1D$ mixing [2]. Similar to the case of $\psi(3686)$ and $\psi(3770)$, the relationship in a $3S$-$2D$ mixing scheme is expressed via the following mixing scheme
@@ II $3S$-$2D$ Mixing scheme for $\psi(3^{3}S_{1})$ and $\psi(2^{3}D_{1})$
where $\theta$ is the mixing angle111 Two commonly used mixing conventions exist for the $\psi(3^{3}S_{1})$ and $\psi(2^{3}D_{1})$ charmonium states. In Refs. [4, 5, 11, 10], the physical states are expressed using a mixing angle $\phi$: $\begin{pmatrix}|\psi^{\prime}_{3S\text{-}2D}\rangle\\ |\psi^{\prime\prime}_{3S\text{-}2D}\rangle\end{pmatrix}=\begin{pmatrix}\cos\phi&\sin\phi\\ -\sin\phi&\cos\phi\end{pmatrix}\begin{pmatrix}|\psi(3^{3}S_{1})\rangle\\ |\psi(2^{3}D_{1})\rangle\end{pmatrix}.$ (2) In contrast, other studies (e.g., Refs. [6, 9]) employ the convention described in Eq. (1). This work adopts the latter convention. Note that the mixing angle $-\phi$ from Refs. [4, 5, 11, 10] is equivalent to our $\theta$ (i.e., $-\phi=\theta$). . Our convention is consistent with Refs. [6, 9]. The $\psi^{\prime}_{3S\text{-}2D}$ and $\psi^{\prime\prime}_{3S\text{-}2D}$ correspond to the charmonium states $\psi(4040)$ and $\psi(4160)$, respectively.
@@
Here, $\Phi_{LS}(3^{3}S_{1})$ and $\Phi_{LS}(2^{3}D_{1})$ represent the $|LSJ\rangle$ basis for the $3S$-wave and $2D$-wave charmonium states, respectively. To quantify the $3S$-$2D$ mixing angle, we employ the Godfrey-Isgur (GI) model, with details provided in Ref. [21]. The Hamiltonian is expressed as
@@
respectively. In this work, we consider $DD$, $DD^{*}$, $D^{*}D^{*}$, $D_{s}D_{s}$, and $D_{s}D_{s}^{*}$ as the primary coupled channels for $\psi(3^{3}S_{1})$ and $\psi(2^{3}D_{1})$, as their respective thresholds lie below the bare mass of $\psi(3^{3}S_{1})$. The transition amplitude $\mathcal{M}_{JL}^{\psi_{S(D)}}=\langle\psi_{(S(D))}|H_{I}|BC\rangle$ describes the interaction between the bare state and the virtual hadronic channels. We adopt the quark pair creation (QPC) model to calculate $\mathcal{M}_{JL}^{\psi_{S(D)}}$ [22, 23]. The subtraction point $M_{J/\psi}$ is taken as the mass of the $J/\psi$ meson in the charmonium system. Previous studies using this method have successfully reproduced reasonable meson masses, as demonstrated in Refs. [24, 25, 26, 27, 3, 28].
In order to obtain the physical mass, we should to solve the eigenvalue equation related to $M$ in Eq. (II). For a general $n\times n$ Hermitian matrix, on could obtain $n$ eigenvalues and $n$ orthogonal eigenvectors. However, in Eq. (II), the Hermitian matrix
@@
naturally contains eigenvalue $M$, which appears in $\Delta M_{S/D}(M)$. In this scheme, it is non-linear eigenvalue problem. For a given $M$, one could obtain two eigenvalues from Eq. (10), but the result which different from the given $M$ may do not satisfy the relationship in Eq. (II). In this scheme, with different physical solutions, the matrix in Eq. (10) may have different matrix elements. Thus the physical eigenvectors may be not orthogonal, which reflect to the solutions with different mixing angles.
@@
Before implementing the $S$-$D$ mixing scheme, we quantitatively analyze the coupling strength of $\psi(3^{3}S_{1})$ and $\psi(2^{3}D_{1})$ to the virtual coupled-channels. For $\psi(3^{3}S_{1})$ and $\psi(2^{3}D_{1})$, the coupled-channel equation Eq. (II) can be decomposed as
@@
When we use the parameter from Ref. [25], the bare masses of $\psi(3^{3}S_{1})$ and $\psi(2^{3}D_{1})$ are estimated to be 4089 MeV and 4172 MeV, respectively. According to our results in Table 1, when the coupled channel effects are taken into account , the bare mass of $\psi(3^{3}S_{1})$ will shift by 81.8 MeV. The $\psi(3^{3}S_{1})$ state predominantly couples to the $D^{*}D^{*}$ channel via the $P$ partial waves, while the $F$ waves provide negligible contributions. The $DD$ ($D_{s}D_{s}$) and $DD^{*}$($D_{s}D_{s}^{*}$) channels provide relatively small contributions to the mass shift of $\psi(3^{3}S_{1})$, mainly through the $P$ waves, respectively.
@@
where $\mathbf{P}$ is the momentum of the final-state mesons, and $M$ is the physical mass.
From Table 1, the open-charm decay width of the $\psi(3^{3}S_{1})$ is predicted to be 45.1 MeV, with the main decay channel being $DD^{*}$. The predicted width of $\psi(3^{3}S_{1})$ is 36.4 MeV and its dominantly decays into $DD$, $DD^{*}$, and $D^{*}D^{*}$ channels.
Table 2: The mass shifts ($\Delta M_{i}$) and decay widths ($\Gamma_{i}$) of the $\psi(3^{3}S_{1})$ and $\psi(2^{3}D_{1})$ states. The values are given in units of MeV. [^1] In this work, $DD$ is used as a shorthand notation to represent the sum of the $D\bar{D}$ and $\bar{D}D$ decay channels, and similar notations are adopted for other decay modes.
@@
Table 3: The $3S$-$2D$ mixing angles for $\psi^{\prime}_{3S\text{-}2D}$ and $\psi^{\prime\prime}_{3S\text{-}2D}$ are induced by the coupled channel effects. The mixing states are expressed as $C_{S1(2)}|\psi(3^{3}S_{1})\rangle$ +$C_{D1(2)}|\psi(2^{3}D_{1})\rangle$.
@@
$\psi^{\prime}_{3S\text{-}2D}$ $-0.992$ $0.128$ $\theta_{1}=7^{\circ}$
@@
By solving Eq. (II) within our established framework, we obtain two solutions for the physical masses of the mixing states $\psi^{\prime}_{3S\text{-}2D}$ and $\psi^{\prime\prime}_{3S\text{-}2D}$, which are $4006.4$ MeV and $4089.5$ MeV, respectively. Additionally, in Table 3, we present the corresponding eigenvectors for these states. Based on the obtained coefficients $C_{S1(2)}$ and $C_{D1(2)}$, the mixing angles are extracted using the relations $\theta_{1}=\arctan(-C_{D1}/C_{S1})$ and $\theta_{2}=\arctan(C_{S2}/C_{D2})$ for $\psi^{\prime}_{3S\text{-}2D}$ and $\psi^{\prime\prime}_{3S-2D}$, respectively [30]. The eigenvectors $(C_{S1(2)},C_{D1(2)})$ represent the mixing coefficients for the $3^{3}S_{1}$ and $2^{3}D_{1}$ components in each physical state. Then, the mixing angles are determined to be
$\theta_{1}=7^{\circ}\quad{\text{for}}\quad\psi^{\prime}_{3S\text{-}2D},$ (14)
@@
$\theta_{2}=10^{\circ}\quad{\text{for}}\quad\psi^{\prime\prime}_{3S\text{-}2D}.$ (15)

### v3

# Is the 3S3S-$2D$ mixing strong for the charmonia $\psi(4040)$ and $\psi(4160)$?

@@ Abstract
In this work, we revisit the 3S3S-$2D$ mixing scheme for the charmonia $\psi(4040)$ and $\psi(4160)$. We introduce a coupled-channel mechanism-distinct from the tensor-force contribution in potential models, which alone is insufficient to induce significant mixing-to describe the mixing between these states. Our analysis yields mixing angles of $\theta_{1}=7^{\circ}$ and $\theta_{2}=10^{\circ}$, inconsistent with the larger angle inferred from experimental data, such as the dilectronic widths of the $\psi(4040)$ and $\psi(4160)$. We discuss possible origins of this discrepancy and emphasize the need for future experiments to resolve it. Precise measurements of the resonance parameters and dilectronic decay widths, via both inclusive and exclusive processes, will be crucial in clarifying this issue.
@@ I Introduction
Although 50 years have passed since the discovery of the $J/\psi$, the properties of charmonium states above the $D\bar{D}$ threshold remain poorly understood, presenting a more complex picture due to insufficient experimental data and a lack of comprehensive theoretical understanding. Among these higher-mass states, the $\psi(4040)$ and $\psi(4160)$ were first observed in $e^{+}e^{-}$ annihilation by the DASP Collaboration in 1978 [1]. Both states are now well-established charmonium resonances and are listed in the Review of Particle Physics by the Particle Data Group (PDG). This inclusion might suggest that their properties are sufficiently understood-yet in reality, the situation remains unclear, as we will discuss below.
$S$-$D$ mixing is a universal phenomenon in charmonium physics. A classic example is the 2S2S-$1D$ mixing scheme proposed by Rosner for $\psi(3770)$ and $\psi(3686)$ [2], which successfully explains the relevant experimental data. Recently, the Lanzhou group suggested 4S4S-$3D$ mixing as a mechanism to interpret the charmoniumlike states $Y(4220)$ and $Y(4380)$ [3]. Similarly, for the well-established charmonia $\psi(4040)$ and $\psi(4160)$, a 3S3S-$2D$ mixing scheme has also been proposed, as discussed in Refs. [4, 5, 6, 7, 8, 9, 10, 11]. The mixing angles can be roughly obtained by fitting either the measured decay widths of the relevant states or the ratio between them. In Table 1, we summarize the current understanding of 3S3S-$2D$ mixing, including the inferred mixing angles and the corresponding approaches employed to extract them.
Table 1: The obtained 3S3S-$2D$ mixing angle for the $\psi(4040)$ and $\psi(4160)$.
@@
Although various methods are employed to determine the mixing angles, their results indicate that a significant 3S3S-$2D$ mixing effect exists. Understanding the nature of the mixing angle is essential, as it has an influence on the assignment of higher vector charmonium states, their properties of mass, production, and decay. Therefore, given the significant 3S3S-$2D$ mixing effect observed in the charmonium states $\psi(4040)$ and $\psi(4160)$, it becomes crucial to identify the underlying dynamical mechanisms responsible for generating such strong mixing.
In this work, we focus on this issue. In the traditional potential model, the tensor term in the effective potential contributes to $S$-$D$ mixing in charmonium, but its effect is too weak to produce a large mixing angle. This necessitates exploring alternative mechanisms. With advances in understanding unquenched effects in hadron spectroscopy, higher-mass states can no longer be accurately described using the quenched approximation. Among these effects, the coupled-channel mechanism has emerged as particularly significant, signaling that hadron spectroscopy is entering a high-precision era [15, 16, 17, 18, 19]. Leveraging these developments, we investigate whether coupled-channel effects could explain the large 3S3S-$2D$ mixing angle reported in earlier studies. However, we find a notable discrepancy between our calculations and prior results: Our analysis yields mixing angles of $\theta_{1}=7^{\circ}$ and $\theta_{2}=10^{\circ}$, differing substantially from previous values.
This discrepancy cannot be overlooked. To resolve it, we scrutinize earlier methodologies and identify a key sensitivity: The mixing angle depends critically on the measured dielectronic widths of the $\psi(4040)$ and $\psi(4160)$. Unfortunately, existing experimental data are insufficient for a precise extraction of the mixing angle. Our findings underscore the urgent need for higher-precision measurements to clarify the $S$-$D$ mixing problem in the $\psi(4040)$ and $\psi(4160)$.
@@
II 3S3S-$2D$ Mixing scheme for $\psi(3^{3}S_{1})$ and $\psi(2^{3}D_{1})$
As mentioned above, charmonia $\psi(3770)$ and $\psi(3686)$ can be well understood by incorporating 2S2S-$1D$ mixing [2]. Similar to the case of $\psi(3686)$ and $\psi(3770)$, the relationship in a 3S3S-$2D$ mixing scheme is expressed via the following mixing scheme
@@
where $\theta$ is the mixing angle [^1] Two commonly used mixing conventions exist for the $\psi(3^{3}S_{1})$ and $\psi(2^{3}D_{1})$ charmonium states. In Refs. [4, 5, 11, 10], the physical states are expressed using a mixing angle $\phi$: $\begin{pmatrix}|\psi^{\prime}_{3S\text{-}2D}\rangle\\ |\psi^{\prime\prime}_{3S\text{-}2D}\rangle\end{pmatrix}=\begin{pmatrix}\cos\phi&\sin\phi\\ -\sin\phi&\cos\phi\end{pmatrix}\begin{pmatrix}|\psi(3^{3}S_{1})\rangle\\ |\psi(2^{3}D_{1})\rangle\end{pmatrix}$. (2) In contrast, other studies (e.g., Refs. [6, 9]) employ the convention described in Eq. (1). This work adopts the latter convention. Note that the mixing angle $\phi$ from Refs. [4, 5, 11, 10] is equivalent to our $\theta$ (i.e., $\phi=\theta$). . Our convention is consistent with Refs. [6, 9]. The $\psi^{\prime}_{3S\text{-}2D}$ and $\psi^{\prime\prime}_{3S\text{-}2D}$ correspond to the charmonium states $\psi(4040)$ and $\psi(4160)$, respectively.
@@
Here, $\Phi_{LS}(3^{3}S_{1})$ and $\Phi_{LS}(2^{3}D_{1})$ represent the $|LSJ\rangle$ basis for the $3S$-wave and $2D$-wave charmonium states, respectively. To quantify the 3S3S-$2D$ mixing angle, we employ the Godfrey-Isgur (GI) model, with details provided in Ref. [21]. The Hamiltonian is expressed as
@@
respectively. In this work, we consider $DD$, $^{*}, DDD^{*}D^{*}$, $D_{s}D_{s}$, and $D_{s}D_{s}^{*}$ as the primary coupled channels for $\psi(3^{3}S_{1})$ and $\psi(2^{3}D_{1})$, as their respective thresholds lie below the bare mass of $\psi(3^{3}S_{1})$. The transition amplitude $\mathcal{M}_{JL}^{\psi_{S(D)}}=\langle\psi_{(S(D))}|H_{I}|BC\rangle$ describes the interaction between the bare state and the virtual hadronic channels. We adopt the quark pair-creation (QPC) model to calculate $\mathcal{M}_{JL}^{\psi_{S(D)}}$ [22, 23]. The subtraction point $M_{J/\psi}$ is taken as the mass of the $J/\psi$ meson in the charmonium system. Previous studies using this method have successfully reproduced reasonable meson masses, as demonstrated in Refs. [24, 25, 26, 27, 3, 28].
In order to obtain the physical mass, we should to solve the eigenvalue equation related to $M$ in Eq. (II). For a general $n\times n$ Hermitian matrix, one could obtain $n$ eigenvalues and $n$ orthogonal eigenvectors. However, in Eq. (II), the Hermitian matrix
@@
naturally contains eigenvalue $M$, which appears in $\Delta M_{S/D}(M)$. In this scheme, it is a nonlinear eigenvalue problem. For a given $M$, one could obtain two eigenvalues from Eq. (10), but the result which is different from the given $M$ may not satisfy the relationship in Eq. (II). In this scheme, with different physical solutions, the matrix in Eq. (10) may have different matrix elements. Thus the physical eigenvectors may be not orthogonal, which reflect to the solutions with different mixing angles.
@@
Before implementing the $S$-$D$ mixing scheme, we quantitatively analyze the coupling strength of $\psi(3^{3}S_{1})$ and $\psi(2^{3}D_{1})$ to the virtual coupled channels. For $\psi(3^{3}S_{1})$ and $\psi(2^{3}D_{1})$, the coupled-channel equation (II) can be decomposed as
@@
When we use the parameter from Ref. [25], the bare masses of $\psi(3^{3}S_{1})$ and $\psi(2^{3}D_{1})$ are estimated to be 4089 and 4172 MeV, respectively. According to our results in Table 2, when the coupled-channel effects are taken into account , the bare mass of $\psi(3^{3}S_{1})$ will shift by 81.8 MeV. The $\psi(3^{3}S_{1})$ state predominantly couples to the $D^{*}D^{*}$ channel via the $P$ partial waves, while the $F$ waves provide negligible contributions. The DDDD ($D_{s}D_{s}$) and $DD^{*}$($D_{s}D_{s}^{*}$) channels provide relatively small contributions to the mass shift of $\psi(3^{3}S_{1})$, mainly through the $P$ waves, respectively.
@@
where $\mathbf{P}$ is the momentum of the final-state mesons and $M$ is the physical mass.
From Table 2, the open-charm decay width of the $\psi(3^{3}S_{1})$ is predicted to be 45.1 MeV, with the main decay channel being $DD^{*}$. The predicted width of $\psi(3^{3}S_{1})$ is 36.4 MeV, and it dominantly decays into $DD$, $DD^{*}$, and $D^{*}D^{*}$ channels.
Table 2: The mass shifts ($\Delta M_{i}$) and decay widths ($\Gamma_{i}$) of the $\psi(3^{3}S_{1})$ and $\psi(2^{3}D_{1})$ states. The values are given in units of MeV.In this work, DDDD is used as a shorthand notation to represent the sum of the $D\bar{D}$ and $\bar{D}D$ decay channels, and similar notations are adopted for other decay modes.
@@
Table 3: The 3S3S-$2D$ mixing angles for $\psi^{\prime}_{3S\text{-}2D}$ and $\psi^{\prime\prime}_{3S\text{-}2D}$ are induced by the coupled-channel effects. The mixing states are expressed as $C_{S1(2)}|\psi(3^{3}S_{1})\rangle$ $C_{D1(2)}|\psi(2^{3}D_{1})\rangle$.
@@
$\psi^{\prime}_{3S\text{-}2D}$ 0.992-0.992 $0.128$ $\theta_{1}=7^{\circ}$
@@
From Eq. (II), one can see that the $\Delta M_{SD}(M)$ term plays a dominant role in determining the mixing angle. The associated uncertainty is mainly governed by the quark pair-creation strength $\gamma$, which simultaneously affects the physical masses, the OZI-allowed two-body strong decay widths, and the mixing angles. To obtain reliable estimations, $\gamma$ is typically taken to be around $0.4$ [25, 27], as fitted from the decay widths of various charmonium states. If $\gamma$ deviates significantly from this value, the resulting physical masses and decay widths become unreasonable. To examine its impact, we consider three representative values $\gamma=0.37$, $0.40$, and $0.43$, and analyze their corresponding effects on $\psi^{\prime}_{3S\text{-}2D}$ and $\psi^{\prime\prime}_{3S\text{-}2D}$.
By solving Eq. (II) within our established framework, we obtain the physical masses of the mixing states $\psi^{\prime}_{3S\text{-}2D}$ and $\psi^{\prime\prime}_{3S\text{-}2D}$, which are $4006.4^{+7.4}_{-8.3}$ and $4089.5^{+8.8}_{-8.6}$ MeV, respectively. Additionally, in Table 3, we present the corresponding eigenvectors for these states. Based on the obtained coefficients $C_{S1(2)}$ and $C_{D1(2)}$, the mixing angles are extracted using the relations $\theta_{1}=\arctan(-C_{D1}/C_{S1})$ and $\theta_{2}=\arctan(C_{S2}/C_{D2})$ for $\psi^{\prime}_{3S\text{-}2D}$ and $\psi^{\prime\prime}_{3S-2D}$, respectively [30]. The eigenvectors $(C_{S1(2)},C_{D1(2)})$ represent the mixing coefficients for the $3^{3}S_{1}$ and $2^{3}D_{1}$ components in each physical state. Then, the mixing angles are determined to be
$\theta_{1}=(7.3^{+0.1}_{-0.4})^{\circ}\quad{\text{for}}\quad\psi^{\prime}_{3S\text{-}2D},$ (14)
@@
$\theta_{2}=(10.4^{+2.0}_{-3.5})^{\circ}\quad{\text{for}}\quad\psi^{\prime\prime}_{3S\text{-}2D}.$ (15)
Based on the above discussion, the mixing angles remain relatively stable with the change of $\gamma$ and are determined to be around $7^{\circ}$ and $10^{\circ}$. For convenience, these two values are adopted in the subsequent numerical calculations.